OpenStudy (anonymous):

a cell supplies a current of 2.2amp through 2ohm resistors when connected in parallel.when same resistors r connected in series it supplies a current of 0.4amp..find emf and internal resistance of the cell

OpenStudy (wasiqss):

how many 2 ohm resistor are there?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Let the internal resistance and emf of the cell be R and E respectively. when resistors are parallel then I=E/Req=E*(1/2+1/2+1/R)=2.2......(1) when resistors are in series then I=E/Req=E/(2+2+R)=0.4............(11) solve (1) and (11) and find the value of E and R.

OpenStudy (anonymous):